A car moving with a speed of 50 km/hr, can be stopped by brakes after at least 6m. If the same car is moving at a speed of 100 km/hr, the minimum stopping distance is ___
1
Expert's answer
2021-03-10T17:20:21-0500

Let's first find the deceleration of the car:


"v^2=v_0^2+2ad,""a=\dfrac{v^2-v_0^2}{2d},""a=\dfrac{0-(50\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s})^2}{2\cdot6\ m}=-16.07\ \dfrac{m}{s^2}."

Finally, we can find the minimum stopping distance if the same car is moving at a speed of 100 km/hr:


"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a},""d=\dfrac{0-(100\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s})^2}{2\cdot(-16.07\ \dfrac{m}{s^2})}=24\ m."